1/4log(x)-log(100)=2
solve the equation by finding the exact solution

Answer by lwsshak3(6755)

1/4log(x)-log(100)=2
solve the equation by finding the exact solution
**
1/4log(x)-log(100)=2
logx/4-2=2
logx/4=4
logx=16
x=10^16
Check:
(1/4)log(10^16)-log(100)
(1/4)(16)-2=4-2=2